Find the kinetic energy (J) of an electron which has a de Broglie wavelength equal to 605 nm.
Answer #1

\lambda = wavelength = 605 x 10-9 m

m = mass = 9.1 x 10-31 kg

using de broglie formula

\lambda = h/mv

v = h/(m\lambda)

v = (6.6 x 10-34) / (9.1 x 10-31 x 605 x 10-9)

v = 1198.8 m/s

Kinetic energy = (0.5) m v2 = (0.5) (9.1 x 10-31) (1198.8)2 = 6.54 x 10-25 J